OpenStudy (anonymous):

A circular pond of radius 6m is surrounded by a path of width 1m. Calculate the area of the path

OpenStudy (anonymous):

A of path = pi (6^2 - 5^2) = 11pi = 34.6 m

OpenStudy (anonymous):

look at it as 2 circle, so you need to subtract the area of the smallest circle (pond) from the biggest circle (path). if the radius of the pond is 6m, then the radius of the bigger circle is 7m. \[A = \pi.r^2\] \[(\pi.7^2) - (\pi.6^2) = 49\pi - 36\pi = 13\pi \] so, 13pi m²